Compute the z-statistic for each given claim (p), the observed proportion (p^), and the given


sample size (𝑛).



𝑝 = 0.53, 𝑝̂= 0.54, 𝑛 = 32



1
Expert's answer
2022-05-31T11:33:04-0400

"z=\frac{\^p-p}{\sqrt{\frac{p(1-p)}{n}}}"

"z=\frac{0.54-0.53}{\sqrt{\frac{0.53(1-0.53)}{32}}}=0.113"